OpenStudy (anonymous):

Consider the linear function f(x)=4-0.5x. Find and simplify the expression for y=f(x+6).

OpenStudy (anonymous):

wow thats tough

jimthompson5910 (jim_thompson5910):

Replace every `x` with `x+6` \[\Large f(x)=4-0.5x\] \[\Large f(x)=4-0.5(x)\] \[\Large f(\color{red}{x})=4-0.5(\color{red}{x})\] \[\Large f(\color{red}{x+6})=4-0.5(\color{red}{x+6})\] I'll let you simplify